# Discrete Messages Improve Communication Efficiency among Isolated Intelligent Agents

### v2

@@ 4.2 Theoretical Validation and Practical Application
Based on the open-source datasets, we conducted the same experiments as in section 4.1 with our own ocular dataset(Figure 7) to further validate our theory. When experimenting with the VQVAE model on this dataset, codebook size $e_{m}\in\mathbb{R}^{512\times 128}$. The dataset is divided into 5 categories based on symptom types, with varying numbers of images in each category. We performed 5 sets of experiments, each involving communications between agents. First, to demonstrate the feasibility of our proposed multi-token discretization, we conducted experiments following Algorithm 2. The experimental results are shown in Figure 8(Right), which once again validate the effectiveness of our theory.
@@
Subsequently, we proceeded to the second approach, which is based on Algorithm 1.The dataset, which originally consisted of only 2750 images, has been expanded to 5000 images by applying data augmentation techniques. Each class now contains 1000 augmented images. Then, we processed the dataset according to the data preprocessing steps outlined in Algorithm 1, and completed the cross-validation experiments. The cross-validation loss under multi-token discretization is shown in Figure 8(Mid).
@@
Finally, we conducted experiments on the core theoretical aspects based on this dataset. The results, shown in Figure 8(Left), indicate that when the number of discrete tokens reaches 32, the overall discrete interactive communication outperforms continuous interactive communication, the former’s average loss is 7.1% lower than that of the latter. The experimental results on the new dataset provided strong evidence to support our conclusions. Communication between agents who are unfamiliar with each other using multi-token discretized information variables is better than using continuous variables.
@@
Figure 8: Experimental results under the medical scenario dataset. Left: Communication loss on three types of models(AutoEncoder, Vector Quantized Variational Autoencoder, Vector Quantized Variational Autoencoder-32token); Mid: Communication losses of multiple agents under multi-token discretization. Right: Test loss during cross-training of multiple agents under multi-token discretization.
4.3 Research on codebook aspects
@@
The experimental results above prove the core theory of this paper: when unfamiliar agents communicate with each other, the use of multi-token discrete semantics is more effective than that of continuous semantics. The subsequent research will mainly focus on the usage patterns of the codebook when communicating with discrete semantics and how to improve the codebook to enhance communication efficiency. In the research, the MNIST and CIFAR10 datasets are primarily used for exploration. Firstly, We investigated the impact of the size of the latent codebook space on the efficiency of discrete communication. We conducted experiments using a single-token VQVAE model following Algorithm 1, with the number of agents $m$ set to 10. In the experiments, we controlled the experimental variable to be the size of the first dimension of the latent space. The result is shown in Figure 9.
@@
Figure 9: The communication test losses under different latent space sizes. With the increase in the size of the codebook, the loss obtained by the 10 pairs of agents during the cross-validation phase shows a downward trend. Left: MNIST. Right: CIFAR10.
@@
It can be observed that as the codebook space increases, the communication loss between different agents decreases overall. Although there is some fluctuation in the subsequent data for the MNIST dataset, we speculate that this is due to the small dataset size and the large codebook space, which results in the model not fully learning and the codebook not being evenly distributed. Therefore, we have reason to believe that as the codebook space expands, agents can capture more patterns when learning the language, thereby further improving the efficiency of discrete communication and enhancing the performance of discrete learning.
@@
After completing the aforementioned experiments, in order to further investigate this direction in-depth, we conducted a study on the utilization of the codebook and some patterns in the VQVAE model. In the following results, our experiments were not conducted according to the aforementioned algorithm, but rather using a single model trained on the official datasets.
@@
Figure 11 shows the proportion of the number of codewords used in each iteration to the total number, under different numbers of discrete tokens. It can be observed that when the number of discrete tokens is greater than or equal to 8, the codebook is effectively utilized throughout the iterations. This undoubtedly contributes to improving the model’s performance.
@@
Figure 13: The extent of difference in understanding of the same language among ten pairs of agents. Section 3 mentions that the understanding of the same language by different agents is not completely identical. However, as learning progresses, their similarity in language comprehension increases. Left: Experiments on MNIST. Mid: Experiments on CIFAR10. Right: The diagram to show that different codebook becomes more and more similar. That is, the codebook distributions in the figure are becoming increasingly similar.
@@
In section 4.3 of the research content, it is stated that increasing the size of the latent codebook space is beneficial for agents to improve their discrete communication efficiency. The patterns of codebook usage will aid in our future research endeavors. Especially the multi-token discretization mechanism, which has improved the issue of uneven codebook usage and mitigated the ”discretization bottleneck”.
5 Conclusion, Limitation and Future Study
Our experiments show that while the efficacy of communication between agents using single-token discrete semantics can rival that with continuous semantics, multi-token discretization before communication notably enhances the quality of information exchange over continous language. Furthermore, multi-token discretization outperforms single-token in terms of system generalization, revealing two key insights: multi-token communications are more effective than single-token approaches, and in contexts where agents encounter diverse languages, sentence construction with multiple discrete tokens yields optimal communication efficiency among isolated intelligent Agents. Additionally, we have put forth a theoretical foundation that contrasts the use of a VAE model against an AE model, noting that agent communicates through the VQVAE model falter in comparison to those facilitated by the VAE model, specific details will be given in the Appendix. The multi-token discretization method we proposed does not perform as well as the VAE, which may be due to optimization difficulties. Identifying the underlying reasons for this discrepancy will be the focus of our future research.
In summary, the multi-token discretization approach we propose outperforms the original single-token discretization method, and compared to continuous language based on the AE model, linguistic communication using multi-token discretization offers a greater advantage for communication among isolated agents.
6 Conflict of Interest
The authors declared no potential conflict of interest with respect to the research, authorship, or publication of this article.
7 Data Availability Statement
All relevant data are within the manuscript and its Appendix.
Appendix
Here we provide additional details about the experimental setup and additional results.
@@ 7 Data Availability Statement
[…]
2:for each iteration $i$ do
3:     Sample input data $x_{i}$
4:     Randomly sample encoder $e_{i}$, $d_{k}$, $c$($i\neq k$, $c$ is $c_{i}$ or $c_{k}$ )
5:     $h_{i}\Leftarrow e_{i}(x_{i})$
6:     $z_{i},L_{quantify}\Leftarrow c_{i}(h_{i})$
7:     $x_{i}^{{}^{\prime}}\Leftarrow d_{i}(z_{i})$
8:     $L_{i}\Leftarrow L_{i}(x_{i}^{{}^{\prime}},x_{i})+L_{quantify}$
9:     optimize $e_{i}$, $c_{i}$ and $d_{i}$ with respect to $L_{i}$
10:     for each pair of agents $j$ do
11:         Sample validate data $x$
12:         $loss_{j}\Leftarrow d_{j}(c_{j}(e_{j}(x)))$
13:         output $loss_{j}$
14:     end for
15:end for
@@
In Figure 6, we have demonstrated that the multi-token discretization mechanism is more effective in terms of communication between agents compared to the single-token discretization mechanism. Prior to this, we had already conducted some experimental work to prove the feasibility of the multi-token discretization mechanism. Algorithm 2 is the process through which we validate the effectiveness of multi-token discrete communication. We varied the intermediate processing architecture between the speaker and listener and recorded the test loss of $m$ pairs of agents throughout the entire training process, as shown in Figure 14.
@@
The curves in the figure indicate that when using cross-training, multi-token discretization indeed outperforms the single-token approach. Furthermore, the results demonstrate a pattern where increasing the number of tokens leads to better performance and faster learning. AE model still exhibits the fastest learning speed. That is, when paired agents learn a language, those that adopt a continuous semantic approach learn the fastest. However, as shown in Figure 5, when these agents interact with new agents, the outcomes are not as good as those of agents that learned through a discrete semantic approach.
@@
Figure 14: Internal test loss during cross-training of multiple agents under multi-token discretization. Our results indicate that multi-token discretization has improved the learning efficiency of agents for communication languages.
@@
In section 4, we mentioned the configuration of experimental parameters. For the four datasets, we adjusted the batch size or the size of the latent code space accordingly. However, all parameters for experiments within the same dataset must remain consistent. During our experiments, we attempted to use the VAE model instead of the AE model as a baseline, and simulated the learning and communication process between a pair of agents using continuous semantics. Similarly, we used Equation 5 to allocate individual datasets to each pair of agents, and the loss during training of the VAE model is represented by Equation 7.
@@
$\mathcal{L}_{VAE}=\mathcal{L}_{recon}+\beta\mathcal{L}_{KL}=\left\|x-x^{\prime}\right\|_{2}+\frac{\beta}{2}\sum_{j=0}^{J}\left(1+log\left(\sigma_{j}^{2}\right)-\mu_{j}^{2}-\sigma_{j}^{2}\right)$ (7)
@@
Regarding all the experiments on the autoencoder (AE) model, we replaced it with a variational autoencoder (VAE) model and repeated the experiments. When repeating the core experiments of Algorithm 1, we found that the agents learning with discrete variables did not achieve very good results when communicating with each other. That is, the loss from communication was relatively high. In contrast, the agents using continuous semantics for communication showed higher efficiency in their exchanges, with lower communication losses.
To further explore and compare the performance of continuous semantic communication based on the VAE model and discrete semantic communication, we have devised a series of experimental setups (see Figure 15). According to these setups, we conducted experiments, where the first major category of experiments involved the speaker’s output being processed discretely for half of the information and the other half either being processed continuously or masked as zero. The second major category involved one half of the information being processed either discretely or continuously, while the other half was masked as zero.
Figure 15: Different methods used by paired agents during training and learning involve transmitting half of the information through continuous language, transferring the other half through discrete language, or masking one half of the information while delivering the remaining half using either continuous or discrete language. There are two models that simulate continuous language: AE and VAE.
The experimental results indicate that regardless of whether pairs of agents are trained and learned through the aforementioned combined model, or learned with singel model by masking half, under the condition of a lower overlap ratio, the relationship between the three types of cross-validation losses is: $AE>VQVAE>VAE$
However, this pattern is not absolute. Our main evaluation metric is the reconstruction loss between unfamiliar agents, as we mentioned in Section 3, where unfamiliarity indicates that their training datasets are not completely identical. When the proportion of overlap in the datasets is low, the discrete communication approach indeed does not perform as well as the continuous communication approach.
@@
Figure 16 represents some results when we trained using a combination of continuous and discrete methods, and used different validation methods during communication validation. The training method here involves splitting the encoder’s output into two parts: one part goes through the latent variable layer of the VAE model, and the other part goes through the codebook layer of the VQVAE model, and then both parts are integrated into the decoder. During communication validation, we only use continuous, discrete, or a combination of continuous and discrete methods for cross-validation. The loss during the training process can be represented by Equation 8.

### v3

@@ IV-B Theoretical Validation and Practical Application
Based on the open-source datasets, we conducted the same experiments as in section IV-A with Diabetic Retinopathy dataset(Figure 7) to further validate our theory. When experimenting with the VQ-VAE model on this dataset, codebook size $e_{m}\in\mathbb{R}^{512\times 128}$. The dataset is divided into $5$ categories based on symptom types, with varying numbers of images in each category. We performed $5$ sets of experiments, each involving communications between agents.
@@
The dataset, which originally consisted of only $2750$ images, has been expanded to $5000$ images by applying data augmentation techniques. Each class now contains $1000$ augmented images. First, we processed the dataset according to the data preprocessing steps outlined in Algorithm 1, and completed the communication-validation experiments. The communication loss under multi-token discretization is shown in Figure 8(Right).
@@
Then, we conducted experiments on the core theoretical aspects based on this dataset. The results, shown in Figure 8(Left), indicate that when the number of discrete tokens reaches $32$, the overall discrete interactive communication outperforms continuous interactive communication, the former’s average loss is 7.1% lower than that of the latter. The experimental results on the new dataset provided strong evidence to support our conclusions. Communication between agents who are unfamiliar with each other using multi-token discretized information variables is better than using continuous variables.
@@
Figure 8: Experimental results under Diabetic Retinopathy dataset. Left: Communication losses on three types of models(AE, VQ-VAE, VQ-VAE-32token); Right: Communication losses of multiple agents under multi-token discretization.
IV-C Research on codebook aspects
@@
The subsequent research will mainly focus on the usage patterns of the codebook when communicating with discrete semantics and how to improve the codebook to enhance communication efficiency. In the research, the MNIST and CIFAR10 datasets are primarily used for exploration. Firstly, We investigated the impact of the size of the latent codebook space on the efficiency of discrete communication. We conducted experiments using a single-token VQ-VAE model following Algorithm 1, with the number of agents $m$ set to $10$. In the experiments, we controlled the experimental variable to be the size of the first dimension of the latent space. The result is shown in Figure 9.
@@
Figure 9: The communication losses under different latent space sizes. With the increase in the size of the codebook, the loss obtained by $10$ agents during the communication-validation phase shows a downward trend. Left: MNIST. Right: CIFAR10.
@@
Although there is some fluctuation in the subsequent data for the MNIST dataset, we speculate that this is due to the small dataset size and the large codebook space. Therefore, we have reason to believe that as the codebook space expands, agents can capture more patterns when learning the language, thereby further improving the efficiency of discrete communication and enhancing the performance of discrete learning.
@@
In order to further investigate this direction in-depth, we conducted a study on the utilization of the codebook and some patterns in the VQ-VAE model. In the following results, our experiments were not conducted according to the aforementioned algorithm, but rather using a single model trained on the official datasets.
@@
Figure 11 shows the proportion of the number of codewords used in each iteration to the total number, under different numbers of discrete tokens. It can be observed that when the number of discrete tokens is greater than or equal to $8$, the codebook is effectively utilized throughout the iterations.
@@
Figure 13: The extent of difference in understanding of the same language among $10$ agents. As learning progresses, their similarity in language comprehension increases. Left: Experiments on MNIST. Mid: Experiments on CIFAR10. Right: The diagram to show that different codebooks become more and more similar.
@@
In section IV-C of the research content, it is stated that increasing the size of the latent codebook space is beneficial for agents to improve their discrete communication efficiency. The patterns of codebook usage will aid in our future research endeavors. Especially the multi-token discretization mechanism, which has improved the issue of uneven codebook usage and mitigated the discretization bottleneck.
V Conclusion, Limitation and Future Study
Our experiments have shown that communication between agents using single-token discrete semantics can achieve comparable efficacy to that of continuous semantics. However, multi-token discretization before communication significantly enhances the quality of information exchange compared to continuous language. Additionally, multi-token discretization outperforms single-token approaches in terms of system generalization. This suggests that multi-token communications are more effective, especially in contexts where agents encounter diverse languages. Furthermore, we have explored the use of a VAE model versus an AE model for communication, and found that agents’ communication facilitated by the VAE model outperforms VQ-VAE model-based communication. We aim to investigate the underlying reasons for this discrepancy in future research. In summary, the multi-token discretization approach we propose outperforms the original single-token discretization method, and compared to continuous communication based on the AE model, using multi-token discretization offers a greater advantage for communication among isolated agents.
Acknowledgments
We would like to express our deep gratitude to the authors of the original model VQ-VAE. Their model provides valuable foundation and inspiration for our research. In addition, we would like to thank other relevant researchers for providing code. URL: https://github.com/zalandoresearch/pytorch-vq-vae?tab=MIT-1-ov-file.
Appendix A Appendix
The models used in this paper were implemented on the PyTorch $1.12.1$ framework, using PyCharm Community Edition $2023.1$ on the Windows platform. The model training was conducted on a single GeForce RTX $3060$ GPU with 8GB of GPU memory, using the CUDA $12.1$ experimental environment. The operating system used was Windows. Here we provide additional details about the experimental setup and additional results.
@@ Appendix A Appendix
1:Train $m$ agents simultaneously;
2:Initialize encoders $E=\{e_{0},...,e_{m}\}$
3:Initialize quantization layers $H=\{h_{0}\left(\cdot\right),...,h_{m}\left(\cdot\right)\}$
4:Initialize decoders $D=\{d_{0},...,d_{m}\}$
5:for each iteration $i$ do
6:     Sample input data $x$
7:     Randomly sample encoder $e_{i}$, $d_{k}$, $c$($i\neq k$, $c$ is $c_{i}$ or $c_{k}$ )
8:     $z_{q}\Leftarrow e_{i}(x)$
9:     $z_{i},L_{quantify}\Leftarrow h_{i}(z_{q})$
10:     $x^{\prime}\Leftarrow d_{i}(z_{i})$
11:     $L_{i}\Leftarrow MSE(x^{\prime},x)+L_{quantify}$
12:     optimize $e_{i}$, $h_{i}$ and $d_{i}$ with respect to $L_{i}$
13:     for each pair of agents $j$ do
14:         Sample validate data $v$
15:         $loss_{j}\Leftarrow d_{j}(h_{j}(e_{j}(v)))$
16:         output $loss_{j}$
17:     end for
18:end for
@@
The CelebA dataset we used in all experiments was $8000$ images extracted from official sources, and in the experiment of Algorithm 1, we divided them into $8$ categories based on their attributes for use by $8$ agents. In Figure 6, we have demonstrated that the multi-token discretization mechanism is more effective in terms of communication between agents compared to the single-token discretization mechanism. Prior to this, we had already conducted some experimental work to prove the feasibility of the multi-token discretization mechanism. Algorithm 2 is the process through which we validate the effectiveness of multi-token discrete communication. For the dataset used in Algorithm 2, MNIST and CIFAR10 are used directly without any changes, while the CelebA and retinal datasets are divided into two categories: training and validation sets. That is, the agents in the algorithm are exposed to the same training or validation sets. We varied the intermediate processing architecture between the speaker and listener and recorded the test loss of $m$ agents throughout the entire training process, as shown in Figure 14.
@@
The curves in the figure indicate that when using cross-training, multi-token discretization indeed outperforms the single-token approach. Furthermore, the results demonstrate a pattern where increasing the number of tokens leads to better performance and faster learning. That’s why we initially wanted to use multi-token discretization mechanism for agents’ communication. AE model still exhibits the fastest learning speed. That is, when agent learn a language, those that adopt a continuous semantic approach learn the fastest. However, as shown in Figure 5, when these agents interact with new agents, the outcomes are not as good as those of agents that learned through a discrete semantic approach.
@@
Figure 14: Internal test loss during cross-training of multiple agents. Compared to the single token, multi-token discretization has improved the learning speed of agents for communication languages.
Figure 15: Different methods for communication between agents. One method is to perform continuous communication on half of the information and discrete communication on the other half. Another method is to mask half of the information and use the remaining half for continuous or discrete communication. Here, two models are used to simulate continuous communication, namely autoencoder (AE) and variational autoencoder (VAE).
@@
In section IV, we mentioned the configuration of experimental parameters. For the four datasets, we adjusted the batch size or the size of the latent code space accordingly. However, all parameters for experiments within the same dataset must remain consistent. During our experiments, we attempted to use the VAE model instead of the AE model as a baseline, and simulated the learning and communication process between a pair of agents using continuous semantics. Similarly, we used Equation 4 to allocate individual datasets to each agent, and the loss during training of the VAE model is represented by Equation 6.
@@
$\mathcal{L}_{VAE}=\mathcal{L}_{recon}+\beta\mathcal{L}_{KL}\\ =\left\|x-x^{\prime}\right\|_{2}+\frac{\beta}{2}\sum_{j=0}^{J}\left(1+log\left(\sigma_{j}^{2}\right)-\mu_{j}^{2}-\sigma_{j}^{2}\right)$ (6)
@@
Regarding all the experiments on the autoencoder (AE) model, we replaced it with a variational autoencoder (VAE) model and repeated the experiments. When repeating the core experiments of Algorithm 1, we found that the agents learning with discrete variables did not achieve very good results when communicating with each other. That is, the loss from communication was relatively high. In contrast, the agents using continuous semantics for communication showed higher efficiency in their exchanges, with lower communication loss.
To further explore and compare the performance of continuous semantic communication based on the VAE model and discrete semantic communication, we have devised a series of experimental setups (see Figure 15). According to these setups, we conducted experiments, where the first major category of experiments involved the speaker’s output being processed discretely for half of the information and the other half either being processed continuously. The second major category involved one half of the information being processed either discretely or continuously, while the other half was masked as zero. We have conducted extensive experiments on the structure of Figure 15, but still have not yielded good results. The experimental results indicate that regardless of whether agents are trained and learned through the aforementioned combined model, or learned with singel model by masking half, under the condition of a lower overlap ratio, the relationship between the three types of cross-validation losses is: $AE>VQVAE>VAE$
However, this pattern is not absolute. Our main evaluation metric is the reconstruction loss between unfamiliar agents, as we mentioned in Section III, where unfamiliarity indicates that their training datasets are not completely identical. When the proportion of overlap in the datasets is low, the performance of discrete communication methods is indeed inferior to continuous communication methods based on VAE.
@@
Figure 16 represents some results when we trained using a combination of continuous and discrete methods, and used different validation methods during communication validation. The training method here involves splitting the encoder’s output into two parts: one part goes through the latent variable layer of the VAE model, and the other part goes through the codebook layer of the VQ-VAE model, and then both parts are integrated into the decoder. During communication validation, we only use continuous, discrete, or a combination of continuous and discrete methods for cross-validation. The loss during the training process can be represented by Equation 7.